Your last submission is used for your score 1. O 0/2 points SerPSET9 23.P 002 My Notes Ask Your Teache (a) Calculate the number of electrons in a small, electrically neutral silver pin that has a mass of 7.0 g. Silver has 47 electrons per atom, and its molar mass is 107.87 g/mol. (b) Imagine adding electrons to the pin until the negative charge has the very large value 3.00 mC. How many electrons are added for every 109 electrons already present? Note that 10 is one billion. It might help to rewrite your answer from part (a) as the number of billions of electrons Noed Help? LWaN
Answer #1

Molar mass of silver = 107.87g per 6.023*10^23 atoms (Avogadro mol)

7g of silver contains (7 /107.87) mol = 0.0649 mol

Number of silver atoms = 0.0649 mol x 6.023*10^23 = 3.91*10^22 atoms

Number of electrons, N = 3.91*10 ^22 atoms x 47 .. .. ►N = 1.84*10^24 electrons♦

b)

Electron charge = 1.60*10^-19 C

Number of electrons in 3.0*10^-3 C .. .. n = 3.0*10^-3 C / (1.60^-19 C) = 1.875*10 ^16 electrons required

Number of [10^9 ] electrons = (1.84*10^24♦ / 1.0*10^9) = 1.85*10^15

Number of electrons added per [10^9] = 1.875*10^16 total electrons / 1.85*10^15 ..

..►= 10.2 per [10^9]